# Anosov representations, strongly convex cocompact groups and weak eigenvalue gaps

### v3

@@ Abstract.
We provide characterizations of Anosov representations of word hyperbolic groups into real semisimple Lie groups in terms of the existence of equivariant limit maps on the Gromov boundary, the Cartan property and the uniform gap summation property introduced by Guichard-Guéritaud-Kassel-Wienhard in [22]. We also study representations of finitely generated groups satisfying weak uniform gaps in eigenvalues and establish conditions to be Anosov. As an application, we also obtain a characterization of strongly convex cocompact subgroups of the projective linear group $\mathsf{PGL}_{d}(\mathbb{R})$.
@@ 1. Introduction
Anosov representations of fundamental groups of closed negatively curved Riemannian manifolds were introduced by Labourie [33] in his study of the Hitchin component. Labourie’s definition was later extended by Guichard–Wienhard in [23] for general word hyperbolic groups. Anosov representations have been extensively studied during the last decade by Guichard–Wienhard [23], Kapovich–Leeb–Porti [27, 28, 29], Guéritaud–Guichard–Kassel–Wienhard [22], Bochi–Potrie–Sam-barino [7], Danciger–Guéritaud-Kassel [17], Zimmer [39] and others, and are now are recognized as a higher rank analogue of convex cocompact representations of word hyperbolic groups into simple Lie groups of real rank $1$. Moreover, recently, there have been introduced certain generalizations of classical Anosov representations for relatively hyperbolic groups, we refer to the work of Kapovich–Leeb [26], Zhu [38] and Weisman [40] for more details.
Based on the existing characterizations established in [23, 22, 27, 28, 29, 7, 31], one may define Anosov representations of a hyperbolic group into a semisimple Lie group in terms of the existence of a pair of well-behaved limit maps from the Gromov boundary of the domain group to the corresponding flag spaces, or entirely in terms of uniform gaps in the Cartan or Lyapunov projection of the image of the representation. The purpose of the present paper is to provide new characterizations and strengthen some of the existing ones. Our characterizations are in terms of the existence of limit maps, the Cartan property (see subsection 1.1) and the uniform gap summation property introduced in [22]. As an application of our main results we also obtain characterizations of strongly convex cocompact subgroups of the projective linear group $\mathsf{PGL}_{d}(\mathbb{R})$ (see subsection 1.3). More generally, we study linear representations of finitely generated groups satisfying weak uniform gaps in eigenvalues and we establish sufficient conditions for the domain group to be word hyperbolic and the representation to be Anosov (see sub-section 1.2). In order to provide such conditions, we study the relation between strong property (U), introduced by Kassel–Potrie in [31], and the uniform gap summation property. More precisely, we prove that a finitely generated non-virtually nilpotent group $\Gamma$ which admits a linear representation with the uniform gap summation property (see Definition 4.7), then $\Gamma$ satisfies strong property (U) which is a condition relating the word length and the stable translation length of certain group elements (see Theorem 1.7).
@@ 1.1. Characterizations in terms of limit maps and the Cartan property
Let $\Gamma$ be an infinite word hyperbolic group, $G$ be a linear, non-compact semisimple Lie group with finitely many connected components and fix $K$ a maximal compact subgroup of $G$. We also fix a Cartan subspace $\mathfrak{a}$ of $\mathfrak{g}$, $\overline{\mathfrak{a}}^{+}$ a closed Weyl chamber of $\mathfrak{a}$, a Cartan decomposition $G=K\exp(\overline{\mathfrak{a}}^{+})K$ and consider the Cartan projection $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$. For a linear form $\varphi\in\mathfrak{a}^{\ast}$ and $H\in\mathfrak{a}$, define $\langle\varphi,H\rangle:=\varphi(H)$.
Every subset $\theta\subset\Delta$ of simple restricted roots of $G$ defines a pair of opposite parabolic subgroups $P_{\theta}^{+}$ and $P_{\theta}^{-}$, well defined up to conjugation. Labourie’s dynamical definition of a $P_{\theta}$-Anosov representation $\rho:\Gamma\rightarrow G$ requires the existence of a pair of continuous $\rho$-equivariant maps from the Gromov boundary $\partial_{\infty}\Gamma$ to the flag spaces $G/P_{\theta}^{+}$ and $G/P_{\theta}^{-}$ called the Anosov limit maps of $\rho$ (see Definition 2.2). Our first characterization of Anosov representations is based on the existence of a pair of transverse continuous and equivariant limit maps on the Gromov boundary of the groups one of which satisfies the Cartan property:
@@
(i) $\rho$ is $P_{\theta}$-divergent.
(ii) There exists a pair of continuous, $\rho$-equivariant transverse maps
@@
Let us now shortly explain the assumptions of Theorem 1.1. We say that two $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ are transverse, if for any two distinct points $x^{+},x^{-}\in\partial_{\infty}\Gamma$ there is $g\in G$ such that $\xi^{+}(x^{+})=gP_{\theta}^{+}$ and $\xi^{-}(x^{-})=gP_{\theta}^{-}$. The representation $\rho:\Gamma\rightarrow G$ is $P_{\theta}$-divergent if for every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ and $\alpha\in\theta$, the sequence $(\langle\alpha,\mu(\rho(\gamma_{n}))\rangle)_{n\in\mathbb{N}}$ goes to infinity. The map $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ satisfies the Cartan property if for every sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ converging to a point $x\in\partial_{\infty}\Gamma$ in the Gromov boundary, then $\xi^{+}(x)=\lim_{n}k_{n}P_{\theta}^{+}$, where $\rho(\gamma_{n})=k_{n}\exp(\mu(\rho(\gamma_{n})))k_{n}^{\prime}$, $k_{n},k_{n}^{\prime}\in K$, is written in the Cartan decomposition of $G$. Examples of maps with this property are the limit maps of an Anosov representation (see [7] and [22, Thm. 1.3 (4) & 5.3 (4)]). We discuss the Cartan property in more detail in §4, where we prove (see Corollary 4.6) that for a Zariski dense representation $\rho:\Gamma\rightarrow G$ a (necessarily unique if it exists) continuous $\rho$-equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{\pm}$ has to satisfy the Cartan property.
In Theorem 1.1 the assumption that the map $\xi^{+}$ satisfies the Cartan is necessary and cannot be dropped (see Example 10.2). Moreover, we do not assume that the image $\rho(\Gamma)$ contains a $P_{\theta}$-proximal element in $G/P_{\theta}^{\pm}$ or that the pair of maps $(\xi^{+},\xi^{-})$ is compatible at some point $x\in\partial_{\infty}\Gamma$, i.e. the intersection $\textup{Stab}_{G}(\xi^{+}(x))\cap\textup{Stab}_{G}(\xi^{-}(x))$ is a parabolic subgroup of $G$. Under the assumption that both maps $(\xi^{+},\xi^{-})$ satisfy the Cartan property, Theorem 1.1 also follows from [29, Thm 1.7]. We explain how Theorem 1.1 is related to [29, Thm. 1.7], [27, Thm. 5.47] and [22, Thm. 1.3] at the end of this section.
@@
(i) There exist $C,c>1$ such that for every $\gamma\in\Gamma$ and $\alpha\in\theta$,
$\big{\langle}\alpha,\mu(\rho(\gamma))\big{\rangle}\geqslant c\log|\gamma|_{\Gamma}-C.$
(ii) There exist $B,b>0$ such that for every $\gamma\in\Gamma$,
$\max_{\alpha\in\theta}\big{\langle}\omega_{\alpha},2\mu(\rho(\gamma))-\mu(\rho(\gamma^{2}))\big{\rangle}\leqslant B\big{(}2|\gamma|_{\Gamma}-|\gamma^{2}|_{\Gamma}\big{)}+b.$
Now let $\rho:\Gamma\rightarrow G$ be a Zariski dense representation which admits a pair of $\rho$-equivariant limit maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$. In [23, Thm. 5.11], Guichard–Wienhard proved that $\rho$ is $P_{\theta}$-Anosov if and only if $\xi^{+}$ and $\xi^{-}$ are compatible and transverse. By using Theorem 1.1 and Corollary 4.6, we obtain the following slightly improved version of their theorem. For a quasi-convex subgroup $H$ of $\Gamma$ we denote by $\iota_{H}:\partial_{\infty}H\xhookrightarrow{}\partial_{\infty}\Gamma$ the Cannon–Thurston map extending the inclusion $H\xhookrightarrow{}\Gamma$.
@@
Let $\Gamma$ be a word hyperbolic group, $H$ a quasiconvex subgroup of $\Gamma$, $G$ a semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and $\rho:\Gamma\rightarrow G$ a Zariski dense representation. Suppose that $\rho$ admits continuous $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$. Then the restriction $\rho|_{H}:H\rightarrow G$ is $P_{\theta}$-Anosov if and only if the maps $\xi^{+}\circ\iota_{H}:\partial_{\infty}H\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}\circ\iota_{H}:\partial_{\infty}H\rightarrow G/P_{\theta}^{-}$ are transverse.
For a matrix $g\in\mathsf{GL}_{d}(\mathbb{R})$ we denote by $\ell_{1}(g)\geq\ldots\geq\ell_{d}(g)$ and $\sigma_{1}(g)\geq\ldots\geq\sigma_{d}(g)$ the eigenvalues and the singular values of $g$ in non-increasing order. Let $\rho_{i}:\Gamma\rightarrow\mathsf{SL}_{m_{i}}(\mathbb{R})$, $i\in\{1,2\},$ be two representations such that $\rho_{2}$ is $P_{1}$-Anosov. We recall that the stretch factors associated with the representations $\rho_{1}$ and $\rho_{2}$ are:
@@
(i) $\rho_{1}$ is $P_{1}$-Anosov.
(ii) $\rho_{1}(\Gamma)$ is contained in a semisimple $P_{1}$-proximal Lie subgroup of $\mathsf{SL}_{m_{1}}(\mathbb{R})$ of real rank $1$.
Then for every $\delta>0$ and $p,q\in\mathbb{N}$ with $\textup{dil}_{-}(\rho_{1},\rho_{2})\leqslant\frac{p}{q}\leqslant\textup{dil}_{+}(\rho_{1},\rho_{2})$, there exists an infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ such that for every $n\in\mathbb{N}$:
${\left|\frac{p}{q}-\frac{\log\sigma_{1}(\rho_{1}(\gamma_{n}))}{\log\sigma_{1}(\rho_{2}(\gamma_{n}))}\right|\leqslant\frac{\delta}{q}\frac{\log|\gamma_{n}|_{\Gamma}}{|\gamma_{n}|_{\Gamma}}}.$
@@
Let $\Gamma$ be a finitely generated group, $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ a representation and fix $11\leqslant i\leqslant d-$ 1. The representation $\rho$ has a weak uniform i-gap in eigenvalues if there exists $\varepsilon>0$ such that for every $\gamma\in\Gamma$ we have
@@
The existence of a uniform $i$-gap in eigenvalues for $\rho$ is not a sufficient condition to guarantee that the representation is Anosov, and it is a natural question to determine additional conditions guaranteeing that this happens. Guéritaud–Guichard–Kassel–Wienhard proved that if $\Gamma$ is word hyperbolic, $\rho$ has a weak uniform $i$-gap in eigenvalues and admits a pair of continuous, $\rho$-equivariant, dynamics preserving and transverse maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{i}(\mathbb{R}^{d})$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{d-i}(\mathbb{R}^{d})$, then $\rho$ is $P_{i}$-Anosov (see [22, Thm. 1.7 (c)]). Kassel–Potrie proved [31, Prop. 4.12] that if $\Gamma$ satisfies weak property (U) (see Definition 5.1) and $\rho$ has a weak uniform $i$-gap in eigenvalues, then $\rho$ has a strong $i$-gap in singular values: there exist $C,c>0$ such that for every $\gamma\in\Gamma$,
@@
hence $\Gamma$ is hyperbolic and $\rho$ is $P_{i}$-Anosov by the work of Kapovich–Leeb–Porti [28] and Bochi–Potrie–Sambarino [7]. The following theorem, motivated by [31, Question 4.9], provides further conditions under which a linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ of a finitely generated group $\Gamma$ with a weak uniform $i$-gap in eigenvalues is $P_{i}$-Anosov and $\Gamma$ is hyperbolic. For the definition of the Floyd boundary we refer the reader to [19], see also §2.
@@
Let $\Gamma$ be a finitely generated infinite group which is not virtually cyclic and fix $|\cdot|_{\Gamma}:\Gamma\rightarrow\mathbb{N}$ a word length function on $\Gamma$. Suppose that $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is a representation which has a weak uniform $i$-gap in eigenvalues for some $11\leq i\leq d-$ 1. Then the following conditions for $\Gamma$ and $\rho$ are equivalent:
(i) $\Gamma$ is word hyperbolic and $\rho$ is $P_{i}$-Anosov.
(ii) There exists a Floyd function $f$ such that the Floyd boundary $\partial_{f}\Gamma$ of $\Gamma$ is uncountable.
(iii) $\Gamma$ admits a representation $\rho_{1}:\Gamma\rightarrow\mathsf{GL}_{m}(\mathbb{R})$ satisfying the uniform gap summation property.
(iv) $\Gamma$ admits a semisimple representation $\rho_{2}:\Gamma\rightarrow\mathsf{GL}_{r}(\mathbb{R})$ with the property
$\lim_{|\gamma|_{\Gamma}\rightarrow\infty}\frac{1}{\log|\gamma|_{\Gamma}}\log\frac{\sigma_{1}(\rho(\gamma))}{\sigma_{r}(\rho(\gamma))}=+\infty.$
We prove that each one of the conditions (ii), (iii) and (iv) implies that $\Gamma$ has strong property (U) (see Definition 5.1), so (i) will follow by the eigenvalue gap characterization from [31, Prop. 1.2]. The uniform gap summation property is a summability condition for gaps between singular values, see [22, Def. 5.2] and Definition 4.7 for the precise definitions. For example, condition (iii) of the previous theorem is satisfied when there exist $11\leq j\leq m-$ 1 and $C,c>1$ such that for every $\gamma\in\Gamma$
@@
[…]
@@
Let $\Gamma$ be a finitely generated group and fix $|\cdot|_{\Gamma}:\Gamma\rightarrow\mathbb{N}$ a word length function on $\Gamma$. Suppose that there exists a Floyd function $f:\mathbb{N}\rightarrow(0,\infty)$ such that the Floyd boundary $\partial_{f}\Gamma$ of $\Gamma$ is non-trivial. Let $H$ be a torsion free subgroup of $\Gamma$ whose limit set $\Lambda(H)$ in $\partial_{f}\Gamma$ contains at least three points. Then there exists a finite subset $F$ of $H$ and $C>0$, depending only on the group $H$, with the property: for every $\gamma\in H$ there exists $g\in F$ such that
$\big{|}g\gamma\big{|}_{\Gamma}-\big{|}g\gamma\big{|}_{\infty}\leqslant C.$
In particular, if $\Gamma$ is virtually torsion free then it satisfies strong property (U).
@@
Let $\Gamma$ be a finitely generated group which is not virtually nilpotent, $G$ a semisimple Lie group and $\theta\subset\Delta$ a subset of simple restricted roots of $G$. Suppose that there exists a representation $\rho:\Gamma\rightarrow G$ which satisfies the uniform gap summation property with respect to $\theta$ and a Floyd function $f:\mathbb{N}\rightarrow(0,\infty)$. Then the Floyd boundary $\partial_{f}\Gamma$ of $\Gamma$ with respect to $f$ is non-trivial. In particular, $\Gamma$ satisfies strong property (U).

### v4

@@ Abstract.
We provide characterizations of Anosov representations of word hyperbolic groups into real semisimple Lie groups in terms of the existence of equivariant limit maps on the Gromov boundary, the Cartan property and the uniform gap summation property introduced by Guichard–Guéritaud–Kassel–Wienhard in [24]. We also study representations of finitely generated groups satisfying weak uniform gaps in eigenvalues and establish conditions to be Anosov. As an application, we also obtain a characterization of strongly convex cocompact subgroups of the projective linear group $\mathsf{PGL}_{d}(\mathbb{R})$.
@@ 1. Introduction
Anosov representations of fundamental groups of closed negatively curved Riemannian manifolds were introduced by Labourie [36] in his study of the Hitchin component. Labourie’s definition was later extended by Guichard–Wienhard in [25] for general word hyperbolic groups. Anosov representations have been extensively studied during the last decade by Guichard–Wienhard [25], Kapovich–Leeb–Porti [29, 30, 31], Bochi–Potrie–Sambarino [9], Guéritaud–Guichard–Kassel–Wienhard [24], Danciger–Guéritaud–Kassel [19], Zimmer [44] and others, and are now are recognized as a higher rank analogue of convex cocompact representations of word hyperbolic groups into simple Lie groups of real rank $1$. Moreover, recently, there have been introduced certain generalizations of classical Anosov representations for relatively hyperbolic groups and other groups; we refer to the work of Kapovich–Leeb [28], Zhu [43] and Weisman [45] for more details.
Based on the existing characterizations established in [25, 24, 29, 30, 31, 9, 33], one may define Anosov representations of a hyperbolic group into a semisimple Lie group in terms of the existence of a pair of well-behaved limit maps from the Gromov boundary of the domain group to the corresponding flag spaces, or entirely in terms of uniform gaps in the Cartan or Lyapunov projection of the image of the representation. The purpose of the present paper is to provide new characterizations and strengthen some of the existing ones. Our characterizations are in terms of the existence of limit maps, the Cartan property (see subsection 1.1) and the uniform gap summation property introduced in [24]. As an application of our main results, we also obtain characterizations of strongly convex cocompact subgroups of the projective linear group $\mathsf{PGL}_{d}(\mathbb{R})$ (see subsection 1.3). More generally, we study linear representations of finitely generated groups satisfying weak uniform gaps in eigenvalues and we establish sufficient conditions for the domain group to be word hyperbolic and the representation to be Anosov (see sub-section 1.2). In order to provide such conditions, we study the relation between strong property (U), introduced by Kassel–Potrie in [33], and the uniform gap summation property. More precisely, we prove that a finitely generated non-virtually nilpotent group $\Gamma$ which admits a linear representation with the uniform gap summation property (see Definition 4.7), then $\Gamma$ satisfies strong property (U) which is a condition relating the word length and the stable translation length of certain group elements (see Theorem 1.7).
@@ 1.1. Characterizations in terms of limit maps and the Cartan property
Let $\Gamma$ be an infinite word hyperbolic group, $G$ be a linear, non-compact semisimple Lie group with finitely many connected components and fix $K$ a maximal compact subgroup of $G$. We also fix a Cartan subspace $\mathfrak{a}$ of $\mathfrak{g}$, $\overline{\mathfrak{a}}^{+}$ a closed Weyl chamber of $\mathfrak{a}$, a Cartan decomposition $G=K\exp(\overline{\mathfrak{a}}^{+})K$ and consider the Cartan projection $\mu:G\rightarrow\overline{\mathfrak{a}}^{+}$.
Every subset $\theta\subset\Delta$ of simple restricted roots of $G$ defines a pair of opposite parabolic subgroups $P_{\theta}^{+}$ and $P_{\theta}^{-}$, well-defined up to conjugation. Labourie’s dynamical definition of a $P_{\theta}$-Anosov representation $\rho:\Gamma\rightarrow G$ requires the existence of a pair of continuous $\rho$-equivariant maps from the Gromov boundary $\partial_{\infty}\Gamma$ to the flag spaces $G/P_{\theta}^{+}$ and $G/P_{\theta}^{-}$ called the Anosov limit maps of $\rho$ (see Definition 2.2). Our first characterization of Anosov representations is based on the existence of a pair of transverse continuous, equivariant limit maps on the Gromov boundary of the domain group, one of which satisfies the Cartan property:
@@
  1. (i)
    $\rho$ is $P_{\theta}$-divergent.
  2. (ii)
    There exists a pair of continuous, $\rho$-equivariant transverse maps
@@
Let us now briefly explain the assumptions of Theorem 1.1. For a representation $\rho:\Gamma\rightarrow G$ of a hyperbolic group $\Gamma$, two $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$ are transverse, if for any two distinct points $x^{+},x^{-}\in\partial_{\infty}\Gamma$ there is $g\in G$ such that $\xi^{+}(x^{+})=gP_{\theta}^{+}$ and $\xi^{-}(x^{-})=gP_{\theta}^{-}$. The representation $\rho:\Gamma\rightarrow G$ is $P_{\theta}$-divergent if for every infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ and $\alpha\in\theta$, the sequence $\big(\alpha(\mu(\rho(\gamma_{n})))\big)_{n\in\mathbb{N}}$ goes to infinity. The map $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ satisfies the Cartan property if for every sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ converging to a point $x\in\partial_{\infty}\Gamma$ in the Gromov boundary, then $\xi^{+}(x)=\lim_{n}k_{n}P_{\theta}^{+}$, where $\rho(\gamma_{n})=k_{n}\exp(\mu(\rho(\gamma_{n})))k_{n}^{\prime}$, $k_{n},k_{n}^{\prime}\in K$, is written in the Cartan decomposition of $G$. Examples of maps with this property are the limit maps of an Anosov representation (see [9] and [24, Thm. 1.3 (4) & 5.3 (4)]). We discuss the Cartan property in more detail in §4, where we prove (see Corollary 4.6) that for any Zariski dense representation $\rho:\Gamma\rightarrow G$ a (necessarily unique if it exists) continuous $\rho$-equivariant map $\xi:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{\pm}$ has to satisfy the Cartan property.
In Theorem 1.1 the assumption that the map $\xi^{+}$ satisfies the Cartan is necessary and cannot be dropped (see Example 10.2). Moreover, we do not assume that the image $\rho(\Gamma)$ contains a $P_{\theta}$-proximal element in $G/P_{\theta}^{\pm}$ or that the pair of maps $(\xi^{+},\xi^{-})$ is compatible at some point $x\in\partial_{\infty}\Gamma$, i.e. the intersection $\textup{Stab}_{G}(\xi^{+}(x))\cap\textup{Stab}_{G}(\xi^{-}(x))$ is a parabolic subgroup of $G$. Under the assumption that both maps $(\xi^{+},\xi^{-})$ satisfy the Cartan property, Theorem 1.1 also follows from [31, Thm 1.7]. We explain how Theorem 1.1 is related to [31, Thm. 1.7], [29, Thm. 5.47] and [24, Thm. 1.3] at the end of this section.
@@
  1. (i)
    There exist $C,c>1$ such that for every $\gamma\in\Gamma$ non-trivial and $\alpha\in\theta$,
    $\alpha\big(\mu(\rho(\gamma))\big)\geqslant c\log|\gamma|_{\Gamma}-C.$
  2. (ii)
    There exist $B,b>0$ such that for every $\gamma\in\Gamma$ and $\alpha\in\theta$,
    $\omega_{\alpha}\big(2\mu(\rho(\gamma))-\mu(\rho(\gamma^{2}))\big)\leqslant B\big(2|\gamma|_{\Gamma}-|\gamma^{2}|_{\Gamma}\big)+b.$
Now let $\rho:\Gamma\rightarrow G$ be a Zariski dense representation which admits a pair of $\rho$-equivariant, continuous limit maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$. In [25, Thm. 5.11], Guichard–Wienhard proved that $\rho$ is $P_{\theta}$-Anosov if and only if $\xi^{+}$ and $\xi^{-}$ are compatible and transverse. By Theorem 1.1 and Corollary 4.6, we obtain the following slightly improved version of their theorem. For a quasi-convex subgroup $H$ of $\Gamma$ we denote by $\iota_{H}:\partial_{\infty}H\xhookrightarrow{}\partial_{\infty}\Gamma$ the Cannon–Thurston map extending the natural inclusion $H\xhookrightarrow{}\Gamma$.
@@
Let $\Gamma$ be a word hyperbolic group, $H$ a quasiconvex subgroup of $\Gamma$, $G$ a semisimple Lie group, $\theta\subset\Delta$ a subset of simple restricted roots of $G$ and $\rho:\Gamma\rightarrow G$ a Zariski dense representation. Suppose that $\rho$ admits continuous, $\rho$-equivariant maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow G/P_{\theta}^{-}$. Then the restriction $\rho|_{H}:H\rightarrow G$ is $P_{\theta}$-Anosov if and only if the maps $\xi^{+}\circ\iota_{H}:\partial_{\infty}H\rightarrow G/P_{\theta}^{+}$ and $\xi^{-}\circ\iota_{H}:\partial_{\infty}H\rightarrow G/P_{\theta}^{-}$ are transverse.
For a matrix $g\in\mathsf{GL}_{d}(\mathbb{R})$ we denote by $\ell_{1}(g)\geq\cdots\geq\ell_{d}(g)$ and $\sigma_{1}(g)\geq\cdots\geq\sigma_{d}(g)$ the moduli of eigenvalues and the singular values of $g$ respectively in non-increasing order. Let $\rho_{i}:\Gamma\rightarrow\mathsf{SL}_{m_{i}}(\mathbb{R})$, $i\in\{1,2\},$ be two representations such that $\rho_{2}$ is $P_{1}$-Anosov. We recall that the stretch factors associated with the representations $\rho_{1}$ and $\rho_{2}$ of $\Gamma$ are:
@@
  1. (i)
    $\rho_{1}$ is $P_{1}$-Anosov.
  2. (ii)
    $\rho_{1}(\Gamma)$ is contained in a semisimple $P_{1}$-proximal Lie subgroup of $\mathsf{SL}_{m_{1}}(\mathbb{R})$ of real rank $1$.
Then for every $\epsilon>0$ and $p,q\in\mathbb{N}$ with $\textup{dil}_{-}(\rho_{1},\rho_{2})\leqslant\frac{p}{q}\leqslant\textup{dil}_{+}(\rho_{1},\rho_{2})$, there exists an infinite sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of elements of $\Gamma$ such that for every $n\in\mathbb{N}$:
$\Bigg|\frac{p}{q}-\frac{\log\sigma_{1}(\rho_{1}(\gamma_{n}))}{\log\sigma_{1}(\rho_{2}(\gamma_{n}))}\Bigg|\leqslant\frac{\epsilon}{q}\cdot\frac{\log|\gamma_{n}|_{\Gamma}}{|\gamma_{n}|_{\Gamma}}$}.
@@
Let $\Gamma$ be a finitely generated group, $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ a representation and fix $1\leqslant i\leqslant d-1$. The representation $\rho$ has a weak uniform i-gap in eigenvalues if there exists $\varepsilon>0$ such that for every $\gamma\in\Gamma$ we have
@@
The existence of a uniform $i$-gap in eigenvalues for $\rho$ is not a sufficient condition to guarantee that the representation is Anosov, and it is a natural question to determine additional conditions guaranteeing that this happens. Guéritaud–Guichard–Kassel–Wienhard proved that if $\Gamma$ is word hyperbolic, $\rho$ has a weak uniform $i$-gap in eigenvalues and admits a pair of continuous, $\rho$-equivariant, dynamics preserving and transverse maps $\xi^{+}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{i}(\mathbb{R}^{d})$ and $\xi^{-}:\partial_{\infty}\Gamma\rightarrow\mathsf{Gr}_{d-i}(\mathbb{R}^{d})$, then $\rho$ is $P_{i}$-Anosov (see [24, Thm. 1.7 (c)]). Kassel–Potrie proved [33, Prop. 4.12] that if $\Gamma$ satisfies weak property (U) (see Definition 5.1) and $\rho$ has a weak uniform $i$-gap in eigenvalues, then $\rho$ has a strong $i$-gap in singular values: there exist $C,c>0$ such that for every $\gamma\in\Gamma$,
@@
hence $\Gamma$ is hyperbolic and $\rho$ is $P_{i}$-Anosov by the work of Kapovich–Leeb–Porti [30] and Bochi–Potrie–Sambarino [9]. The following theorem, motivated by [33, Ques. 4.9], provides further conditions under which a linear representation $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ of a finitely generated group $\Gamma$ with a weak uniform $i$-gap in eigenvalues is $P_{i}$-Anosov and $\Gamma$ is hyperbolic. For the definition of the Floyd boundary we refer the reader to [21], see also §2.
@@
Let $\Gamma$ be a finitely generated infinite group which is not virtually cyclic and fix $|\cdot|_{\Gamma}:\Gamma\rightarrow\mathbb{N}$ a word length function on $\Gamma$. Suppose that $\rho:\Gamma\rightarrow\mathsf{GL}_{d}(\mathbb{R})$ is a representation which has a weak uniform $i$-gap in eigenvalues for some $1\leq i\leq d-1$. Then the following conditions for $\Gamma$ and $\rho$ are equivalent:
  1. (i)
    $\Gamma$ is word hyperbolic and $\rho$ is $P_{i}$-Anosov.
  2. (ii)
    There exists a Floyd function $f$ such that the Floyd boundary $\partial_{f}\Gamma$ of $\Gamma$ is uncountable.
  3. (iii)
    $\Gamma$ admits a representation $\rho_{1}:\Gamma\rightarrow\mathsf{GL}_{m}(\mathbb{R})$ satisfying the uniform gap summation property.
  4. (iv)
    $\Gamma$ admits a semisimple representation $\rho_{2}:\Gamma\rightarrow\mathsf{GL}_{r}(\mathbb{R})$ with the property
    $\lim_{|\gamma|_{\Gamma}\rightarrow\infty}\frac{\log\sigma_{1}(\rho_{2}(\gamma))-\log\sigma_{r}(\rho_{2}(\gamma))}{\log|\gamma|_{\Gamma}}=+\infty.$
We prove that each one of the conditions (ii), (iii) and (iv) implies that $\Gamma$ has strong property (U) (see Definition 5.1), so (i) will follow by the eigenvalue gap characterization from [33, Prop. 1.2]. The uniform gap summation property is a summability condition for gaps between singular values, see [24, Def. 5.2] and Definition 4.7 for the precise definitions. For example, condition (iii) of the previous theorem is satisfied when there exist $1\leq j\leq m-1$ and $C,c>1$ such that for every $\gamma\in\Gamma$
@@
For the proof of implication $\textup{(ii)}\Rightarrow\textup{(i)}$ in Theorem 1.6 we establish that a torsion-free finitely generated group whose Floyd boundary is uncountable, satisfies strong property (U).
@@
Let $\Gamma$ be a finitely generated group and fix $|\cdot|_{\Gamma}:\Gamma\rightarrow\mathbb{N}$ a word length function on $\Gamma$. Suppose that there exists a Floyd function $f:\mathbb{N}\rightarrow(0,\infty)$ such that the Floyd boundary $\partial_{f}\Gamma$ of $\Gamma$ is non-trivial. Let $H$ be a torsion-free subgroup of $\Gamma$ whose limit set $\Lambda(H)$ in $\partial_{f}\Gamma$ contains at least three points. Then there exists a finite subset $F$ of $H$ and $C>0$, depending only on $H$, with the property: for every $\gamma\in H$ there exists $g\in F$ such that
$|g\gamma|_{\Gamma}-|g\gamma|_{\infty}\leqslant C.$
In particular, if $\Gamma$ is virtually torsion-free then it satisfies strong property (U).
@@
Let $\Gamma$ be a finitely generated group which is not virtually nilpotent, $G$ a semisimple Lie group and $\theta\subset\Delta$ a subset of simple restricted roots of $G$. Let $\rho:\Gamma\rightarrow G$ be a representation which satisfies the uniform gap summation property with respect to $\theta$ and a Floyd function $f:\mathbb{N}\rightarrow(0,\infty)$. Then the Floyd boundary $\partial_{f}\Gamma$ of $\Gamma$ with respect to $f$ is non-trivial. In particular, $\Gamma$ satisfies strong property (U).